# Engineering interaction potentials for stabilizing quantum quasicrystal phases

### v1

@@ I Introduction
The self-assembly of complex phases such as cluster crystals Neuhaus and Likos (2011); Cinti et al. (2014); de Mello et al. (2023); Mendoza-Coto et al. (2021a), quasicrystals Shechtman et al. (1984); Engel and Trebin (2007); Talapin et al. (2009); Fischer et al. (2011); Wasio et al. (2014); Barkan et al. (2014); Dotera et al. (2014); Pupillo et al. (2020); Mendoza-Coto et al. (2022); Grossklags et al. (2024), supersolids Cinti et al. (2010); Zhang et al. (2019); Böttcher et al. (2019); Tanzi et al. (2019); Chomaz et al. (2019); Blakie et al. (2020), and topological phases Berg et al. (2007); Barci et al. (2013); Mendoza-Coto et al. (2015); van Bijnen and Pohl (2015); Mendoza-Coto et al. (2017, 2020); Radzihovsky (2020) is a central topic of many-body physics, with applications in soft matter Prestipino and Saija (2014); Prestipino et al. (2018a); Mendoza-Coto et al. (2021b); Podoliak et al. (2023); Xia and Han (2024) and hard condensed matter Bianconi (2000); Bianconi et al. (2001); Díaz-Méndez et al. (2011); Fernandes et al. (2012); Mendoza-Coto and Stariolo (2012); Fernandes et al. (2014); Mendoza-Coto et al. (2016, 2019). In particular, quasicrystals in two and three dimensions are characterized by a lack of spatial periodicity while exhibiting long-range $n$-fold rotational symmetry (with $n>6$) forbidden by conventional crystallography Levine and Steinhardt (1986); Socolar et al. (1986); Lifshitz (2011). In classical systems of particles, the formation of cluster quasicrystals typically arise in the presence of isotropic soft-core pair potentials with competing length scales Barkan et al. (2014); Dotera et al. (2014), like in the case of dendritic macromolecules Zeng et al. (2004); Dotera (2011). The existence of two degenerate minima in the Fourier transform of the interaction at the appropriate wave vectors controls the self-assembly process of quasicrystals Barkan et al. (2014). In this scenario, the specific values chosen for the ratio between characteristic length scales in the system determines the rotational symmetry of the structure to be stabilized. Such interplay between length scales has been extensively studied in the Lifshitz-Petrich-Gaussian model Barkan et al. (2014); Jiang et al. (2017) due to its mathematical flexibility, which allows the stabilization of periodic and aperiodic phases as the minima structure of the potential and thermal fluctuations are tuned.
@@
These insights have recently guided the search for novel quantum phases, showing that the combined effect of tailored interactions and quantum fluctuations can similarly stabilize quantum quasicrystals Pupillo et al. (2020); Mendoza-Coto et al. (2022); Grossklags et al. (2024); Ciardi et al. (2023). Interestingly, all studies focusing on quantum self-assembled quasicrystals analogous to the one considered by Barkan et al. Barkan et al. (2014), but at zero temperature, have been related to the stabilization of a dodecagonal phase. In this case, the literature Heinonen et al. (2019); Mendoza-Coto et al. (2022) seems to indicate that to enhance stability or even to achieve the stabilization of these phases at all, either the higher momentum minimum should be deeper than the low momentum minimum, or extra resonant incommensurate minima should be added to the pair interaction potential. In this scenario, to push forward the comprehension of the necessary properties of the pair interaction potential for the stabilization of quantum quasicrystals, it is essential to consider scenarios with different rotational symmetries.
@@
Ultracold atoms provide a versatile platform for investigating quantum quasicrystals. With precise control over both internal and external degrees of freedom, these systems have been used to simulate a wide range of quantum phases Gopalakrishnan et al. (2009, 2010); Vaidya et al. (2018); Farokh Mivehvar and Ritsch (2021); Zhang et al. (2021); Karpov and Piazza (2022). However, the realization of quasicrystals Mivehvar et al. (2019a); Farokh Mivehvar and Ritsch (2021) and other exotic phases Guo et al. (2019, 2021) often requires long-range or oscillatory interactions, which are not typically accessible in systems of quantum gases. One novel approach involves the use of a high-finesse optical cavity and a beam shaped laser scanning the system to produce tailored effective sign-changing interactions between the atoms of the system Bonifacio et al. (2024a). In this work, the authors show that the effective interaction having the form $\hat{V}\left(\bm{q}\right)=g_{\text{aa}}-\Delta\leavevmode\nobreak\ \hat{\Omega}\left(\bm{q}\right)^{2}$, is controlled by the Fourier transform of the Rabi frequency profile of the laser $\hat{\Omega}\left(\bm{q}\right)$, the original contact interaction between atoms $g_{\text{aa}}$ and the parameter $\Delta$ accounting for microscopic details of the quantum mechanism involved in the interaction process. The resulting expression offers an unprecedented degree of customization, providing the same level of flexibility of the Lifshitz-Petrich-Gaussian model for the production of classical quasicrystals and other modulated phases. Although this pioneering protocol has not been tested experimentally yet, similar setups have been successfully used already to stabilize supersolids Baumann et al. (2010); Klinder et al. (2015a), spin textured phases Mivehvar et al. (2019b); Guo et al. (2019); Farokh Mivehvar and Ritsch (2021), and extended Hubbard models Klinder et al. (2015b); Landig et al. (2016).
@@
In the present paper, we use a mean field spectral variational approach Mendoza-Coto et al. (2022); Grossklags et al. (2024); Lima et al. (2025) to explore the engineering principles of the pair interaction potential in Fourier space for the stabilization of three prominent quantum quasicrystal phases with octagonal, decagonal, and dodecagonal rotational symmetries. We systematically consider the effects of having two or more competing length scales in the pair potential properly tuned in order to stabilize the corresponding quasicrystal phase and enhance its extension in the ground state phase diagram. We show that, in contrast to the classical scenario, the stabilization of quantum quasicrystals with different rotational symmetries requires pair interaction potentials with a very distinct minima structure. Next, we investigate the reasons for the much greater stability of the dodecagonal structure in comparison to the decagonal and octagonal patterns and present a geometric argument clarifying the origin of such phenomenon. Moreover, to deepen our understanding on the reported phases, we analyze the behavior of the superfluid fraction in three particular cases showing that quasicrystalline order can coexist with global superfluidity, exhibiting a phase analogous to the supersolid phase. We employed two contrasting methods to study the superfluidity, the mean field spectral variational method and the numerical solution of the real-space Gross-Pitaevskii equation in imaginary time, showing that both techniques provide equivalent results. Finally, we discuss the results and present the concluding remarks of our work.
@@ II Model and Method
We consider a two-dimensional gas of $N$ bosonic particles with mass $m$ at zero temperature interacting with a pair potential $V\left(\bm{x}\right)$. Within the mean field approximation, valid at high density of particles and weak enough interactions, the total energy of the system can be expressed as
$\begin{split}E\left[\phi\right]=&\frac{\hbar^{2}}{2m}\int d^{2}{x}\lvert\bm{\nabla}\phi\left(\bm{x}\right)\rvert^{2}+\frac{1}{2}\int d^{2}{x}d^{2}{x}^{\prime}V\left(\bm{x}-\bm{x}^{\prime}\right)\\ &\times\lvert\phi\left(\bm{x}\right)\rvert^{2}\lvert\phi\left(\bm{x}^{\prime}\right)\rvert^{2},\end{split}$ (1)
where $\phi\left(\bm{x}\right)$ stands for the wave function of the condensate, satisfying the normalization condition $\int d^{2}{x}\lvert\phi\left(\bm{x}\right)\rvert^{2}=N$. Moreover, we consider that the pair interaction potential is written in momentum space as $\hat{V}\left(\bm{q}\right)=g_{\text{aa}}-\Delta\leavevmode\nobreak\ \hat{\Omega}\left(\bm{q}\right)^{2}$, such model of pair interaction was presented and discussed earlier in the introduction section. For a study of the stabilization conditions of distinct quasicrystal structures, we consider a rather simple laser beam model in momentum space $\hat{\Omega}\left(\bm{q}\right)$, given by the superposition of several Gaussian profiles centered at different momenta with negligible overlapping between them. As a consequence, the pair interaction potential $\hat{V}\left(\bm{q}\right)$ exhibits a series of local minima properly positioned in order to favor the stabilization of a particular quasicrystalline pattern Mendoza-Coto et al. (2022); Grossklags et al. (2024).
To stabilize a quasicrystal exhibiting $n$-fold rotational symmetry with characteristic momentum $q_{1}$, it is expected that in addition to $q_{1}$, other resonant momenta associated with the corresponding quasicrystal structure should also be excited. In general, if we combine two wave vectors of the extended basis $\bm{q}_{i}=q_{1}\left(\cos\left(2\pi i/n\right),\sin\left(2\pi i/n\right)\right)$ with $i=0,1,\dots,n-1$, we obtain the first group of secondary characteristic wave vectors of the corresponding pattern. For the three simplest and most relevant quasicrystalline structures, octagonal, decagonal, and dodecagonal, $n=8,10$, and $12$, respectively, the first group of secondary wave vectors contains three or more different sets of vectors with an incommensurate length with respect to the characteristic wave vector. Taking this into account, our approach to study the stabilization of quasicrystalline patterns involves separately exciting these length scales in $\hat{\Omega}\left(\bm{q}\right)$, thus clarifying the necessary number of minima in $\hat{V}\left(\bm{q}\right)$ for the desired quasicrystal stabilization. In this way, we consider
@@
$\begin{split}\frac{E\left[\psi\right]}{N\epsilon_{0}}=&\frac{1}{2}\int\frac{d^{2}{r}}{A}\lvert\bm{\nabla}\psi\left(\bm{r}\right)\rvert^{2}+\frac{\lambda^{2}U\rho}{2}\int\frac{d^{2}{r}d^{2}{r}^{\prime}}{A}\\ &v\left(\bm{r}-\bm{r}^{\prime}\right)\lvert\psi\left(\bm{r}\right)\rvert^{2}\lvert\psi\left(\bm{r}^{\prime}\right)\rvert^{2},\end{split}$ (3)
where $A$ stands for the area of the system and $\psi\left(\bm{r}\right)$ represents the normalized wave function, satisfying $\int_{A}d^{2}{r}\lvert\psi\left(\bm{r}\right)\rvert^{2}=A$. Furthermore, the effective pair interaction potential $Uv\left(\bm{r}\right)$ corresponds originally to $V\left(\bm{x}\right)/\epsilon$. Here, we have introduced the dimensionless parameter $U$ to quantify the intensity of the pair interaction potential and $v\left(\bm{r}\right)$ to encode the information about its spatial variation. To avoid any ambiguity in this definition we impose additionally that $\hat{v}(k=1)=-1$, without generality loss. Under these conditions, the local dimensionless density of particles is given by $\rho\left(\bm{r}\right)=\lambda^{2}\rho\lvert\psi\left(\bm{r}\right)\rvert^{2}$, where $\rho$ represents the average particles density. It is worth noticing that within our framework the dimensionless parameter $\lambda^{2}U\rho$ controls the relative intensity of the potential energy contribution in comparison to the kinetic energy in the ground-state energy functional, or in other words the intensity of the quantum fluctuations (zero point motion) in our system. In this way, $\lambda^{2}U\rho$ is naturally one of the running parameter of the phase diagrams for the models considered.
To proceed, we consider that the ground-state wave function $\psi\left(\bm{r}\right)$ minimizing the energy-per-particle functional can be written in a Fourier basis. In this way, we propose Zhang et al. (2019); Mendoza-Coto et al. (2022)
@@
where $c\left(\bm{k}_{j}\right)$ and $\bm{k}_{j}$ represent the Fourier amplitudes and wave vectors of the expansion. The set of wave vectors and amplitudes $\{\bm{k}_{j},c\left(\bm{k}_{j}\right)\}$ defines the type of modulated pattern under consideration and its symmetries. These wave vectors are constructed as all possible integers combinations of a vector basis $\{\bm{k}_{j,0}\}$, the number of elements of the basis defines the rank of the solution. In two dimensions, periodic pattern solutions correspond to configurations of rank 2 or smaller, while quasicrystals have a rank higher than 2. In Table 1, we present the list of solutions considered as well as its wave vector basis, here the variational parameter $k_{m}$ sets the scale of the wave vectors lattice. In terms of possible solutions, we consider all periodic patterns typically found in quasicrystal forming systems Barkan et al. (2014): one dimensional modulations, i.e., stripes, $2$D hexagonal and square crystalline patterns, as well as two different kinds of compressed hexagonal patterns exciting simultaneously the two dominant characteristic wave vectors of the pair potential Lifshitz and Petrich (1997); Jiang et al. (2016); Yin et al. (2021) and finally, $n$-fold rotationally symmetric quasicrystals with $n=\{8,10,12\}$. Within the mean field spectral variational analysis, the set of Fourier amplitudes $\{c_{j}\}$ and the characteristic momentum $k_{m}$ for each possible solution are determined from the minimization of the corresponding energy-per-particle functional Prestipino et al. (2018b, a). Finally, the ground state is calculated selecting the solution with the lowest energy. For the special case of the homogeneous solution, all coefficients $c_{j}$ vanish except for $c_{0}=1$.
Pattern Vector basis $\bm{k}_{j,0}$ Index range
@@
Table 1: Modulated patterns considered in this work and their corresponding wave vector basis $\bm{k}_{j,0}$. The list of of possible solutions includes: stripes, square and hexagonal crystalline patterns, compressed hexagonal patterns exciting simultaneously wave vectors on the ratio $\tilde{k}_{2}$ (see Table 2) and quasicrystalline patterns with eight- ($8$-QC), ten- ($10$-QC) or twelve-fold ($12$-QC) rotational symmetry.

### v2

@@ I Introduction
The self-assembly of complex phases such as cluster crystals [neuhaus2011phonon, cinti2014defect, de2023ultrasoft, mendoza2021ground], quasicrystals [shechtman1984metallic, engel2007self, talapin2009quasicrystalline, fischer2011colloidal, wasio2014self, barkan2014controlled, dotera2014mosaic, pupillo2020quantum, mendoza2022exploring, grossklags2024self, mendoza2024exploring], supersolids [cinti2010supersolid, zhang2019supersolidity, bottcher2019transient, tanzi2019observation, chomaz2019long, blakie2020supersolidity], and topological phases [berg2007dynamical, barci2013nematic, mendoza2015nature, van2015quantum, mendoza2017quantum, mendoza2020topological, radzihovsky2020quantum] is a central topic of many-body physics, with applications in soft matter [prestipino2014hexatic, prestipino2018freezing, mendoza2021cluster, podoliak2023undulations, xia2024simple, mendoza2024melting] and hard condensed matter [bianconi2000superstripes, bianconi2001superstripes, diaz2011dynamics, fernandes2012preemptive, mendoza2012coarse, fernandes2014drives, mendoza2016modulated, mendoza2019mechanism]. In particular, quasicrystals in two and three dimensions are characterized by a lack of spatial periodicity while exhibiting long-range $n$-fold rotational symmetry (with $n>6$) forbidden by conventional crystallography [levine1986quasicrystals, socolar1986phonons, lifshitz2011symmetry]. One singular realization of this kind of system are the so-called cluster quasicrystals [barkan2014controlled, pupillo2020quantum, mendoza2025excitations]. These systems are typically produced by isotropic soft-core pair potentials with a degenerate two minima structure in momentum space properly tuned, at moderate temperatures [zeng2004supramolecular, dotera2011quasicrystals]. To be precise, the relative position of the two competing unstable wave vectors controls the self-assembling process and hence the rotational symmetry of the stabilized quasicrystal [barkan2014controlled]. Such interplay between length scales has been extensively studied using the Lifshitz-Petrich-Gaussian model [barkan2014controlled, jiang2017stability] due to its mathematical flexibility, which allows the stabilization of periodic and aperiodic phases as the minima structure of the potential and thermal fluctuations are tuned.
@@
These insights have recently guided the search for novel quantum phases, showing that the combined effect of tailored interactions and quantum fluctuations can similarly stabilize quantum quasicrystals [pupillo2020quantum, mendoza2022exploring, ciardi2023quasicrystalline, grossklags2024self]. Interestingly, all studies focusing on quantum self-assembled quasicrystals analogous to the one considered by Barkan et al. [barkan2014controlled], but at zero temperature, have been related to the stabilization of a dodecagonal phase. In this case, the literature [heinonen2019quantum, mendoza2022exploring] seems to indicate that to enhance stability or even to achieve the stabilization of these phases at all, either the higher momentum minimum should be deeper than the low momentum minimum, or extra characteristic incommensurate minima should be added to the pair interaction potential. In this scenario, to push forward the comprehension of the necessary properties of the pair interaction potential for the stabilization of quantum quasicrystals, it is essential to consider scenarios with different rotational symmetries.
@@
Ultracold atoms provide a versatile platform for investigating quantum quasicrystals. With precise control over both internal and external degrees of freedom, these systems have been used to simulate a wide range of quantum phases [gopalakrishnan2009emergent, gopalakrishnan2010atom, vaidya2018tunable, mivehvar2021cavity, zhang2021self, karpov2022light]. However, the realization of quasicrystals [mivehvar2019emergent, mivehvar2021cavity] and other exotic phases [guo2019sign, guo2021optical] often requires long-range or oscillatory interactions, which are not typically accessible in systems of quantum gases. One novel approach involves the use of a high-finesse optical cavity and a beam shaped laser scanning the system to produce tailored effective sign-changing interactions between the atoms of the system [bonifacio2024laser]. In this work, the authors show that the effective interaction having the form $\hat{V}\left(\bm{q}\right)=g_{\text{aa}}-\Delta\penalty 10000\ \hat{\Omega}\left(\bm{q}\right)^{2}$, is controlled by the Fourier transform of the Rabi frequency profile of the laser $\hat{\Omega}\left(\bm{q}\right)$, the original contact interaction between atoms $g_{\text{aa}}$ and the parameter $\Delta$ accounting for microscopic details of the quantum mechanism involved in the interaction process. The resulting expression offers an unprecedented degree of customization, providing the same level of flexibility of the Lifshitz-Petrich-Gaussian model for the production of classical quasicrystals and other modulated phases. Although this pioneering protocol has not been tested experimentally yet, similar setups have been successfully used already to stabilize supersolids [baumann2010dicke, klinder2015dynamical], spin textured phases [mivehvar2019cavity, guo2019sign, mivehvar2021cavity], and extended Hubbard models [klinder2015observation, landig2016quantum].
@@
In the present paper, we use a mean field spectral variational approach [mendoza2022exploring, grossklags2024self, lima2025supersolid] to explore the engineering principles of the pair interaction potential in Fourier space for the stabilization of three prominent quantum quasicrystal phases with octagonal, decagonal, and dodecagonal rotational symmetries. We systematically consider the effects of having two or more competing length scales in the pair potential properly tuned in order to stabilize the corresponding quasicrystal phase and enhance its extension in the ground state phase diagram. We show that, in contrast to the classical scenario, the stabilization of quantum quasicrystals with different rotational symmetries requires pair interaction potentials with a very distinct minima structure. Next, we investigate the reasons for the much greater stability of the dodecagonal structure in comparison to the decagonal and octagonal patterns and present a geometric argument clarifying the origin of such phenomenon. Moreover, to deepen our understanding on the reported phases, we analyze the behavior of the superfluid fraction in three particular cases showing that quasicrystalline order can coexist with global superfluidity, exhibiting a phase analogous to the supersolid phase. We employed two contrasting methods to study the superfluidity, the mean field spectral variational method and the numerical solution of the real-space Gross-Pitaevskii equation in imaginary time, showing that both techniques provide equivalent results. Finally, we discuss the results and present the concluding remarks of our work.
@@ II Model and Method
We consider a two-dimensional gas of $N$ bosonic particles with mass $m$ at zero temperature interacting with a pair potential $V(\bm{x})$. Within the mean field approximation, valid at high density of particles and weak enough interactions, the total energy of the system can be expressed as
$\begin{split}E\left[\phi\right]=&\frac{\hbar^{2}}{2m}\int d^{2}{x}\lvert\bm{\nabla}\phi\left(\bm{x}\right)\rvert^{2}+\frac{1}{2}\int d^{2}{x}d^{2}{x}^{\prime}V(\bm{x}-\bm{x}^{\prime})\\ &\times\lvert\phi\left(\bm{x}\right)\rvert^{2}\lvert\phi\left(\bm{x}^{\prime}\right)\rvert^{2},\end{split}$ (1)
where $\phi\left(\bm{x}\right)$ stands for the wave function of the condensate, satisfying the normalization condition $\int d^{2}{x}\lvert\phi\left(\bm{x}\right)\rvert^{2}=N$. Moreover, we consider that the pair interaction potential is written in momentum space as $\hat{V}(\bm{q})=g_{\text{aa}}-\Delta\penalty 10000\ \hat{\Omega}\left(\bm{q}\right)^{2}$, such model of pair interaction was presented and discussed earlier in the introduction section. For a study of the stabilization conditions of distinct quasicrystal structures, we consider a rather simple laser beam model in momentum space $\hat{\Omega}\left(\bm{q}\right)$, given by the superposition of several Gaussian profiles centered at different momenta with negligible overlapping between them. As a consequence, the pair interaction potential $\hat{V}(\bm{q})$ exhibits a series of local minima properly positioned in order to favor the stabilization of a particular quasicrystalline pattern [mendoza2022exploring, grossklags2024self].
To stabilize a quasicrystal exhibiting $n$-fold rotational symmetry with characteristic momentum $q_{1}$, it is expected that in addition to $q_{1}$, other higher order momenta of the corresponding quasicrystal structure should also be excited. In general, if we combine two wave vectors of the extended basis $\bm{q}_{i}=q_{1}\left(\cos\left(2\pi i/n\right),\sin\left(2\pi i/n\right)\right)$ with $i=0,1,\dots,n-1$, we obtain the first group of secondary characteristic wave vectors of the corresponding pattern. For the three simplest and most relevant quasicrystalline structures, octagonal, decagonal, and dodecagonal, $n=8,10$, and $12$, respectively, the first group of secondary wave vectors contains three or more different sets of vectors with an incommensurate length with respect to the characteristic wave vector. Taking this into account, our approach to study the stabilization of quasicrystalline patterns involves separately exciting these length scales in $\hat{\Omega}\left(\bm{q}\right)$, thus clarifying the necessary number of minima in $\hat{V}(\bm{q})$ for the desired quasicrystal stabilization. In this way, we consider
@@
$\begin{split}\frac{E\left[\psi\right]}{N\epsilon_{0}}=&\frac{1}{2}\int\frac{d^{2}{r}}{A}\lvert\bm{\nabla}\psi\left(\bm{r}\right)\rvert^{2}+\frac{\lambda^{2}U\rho}{2}\int\frac{d^{2}{r}d^{2}{r}^{\prime}}{A}\\ &v(\bm{r}-\bm{r}^{\prime})\lvert\psi\left(\bm{r}\right)\rvert^{2}\lvert\psi\left(\bm{r}^{\prime}\right)\rvert^{2},\end{split}$ (3)
where $A$ stands for the area of the system and $\psi\left(\bm{r}\right)$ represents the normalized wave function, satisfying $\int_{A}d^{2}{r}\lvert\psi\left(\bm{r}\right)\rvert^{2}=A$. Furthermore, the effective pair interaction potential $Uv(\bm{r})$ corresponds originally to $V(\bm{x})/\epsilon$. Here, we have introduced the dimensionless parameter $U$ to quantify the intensity of the pair interaction potential and $v(\bm{r})$ to encode the information about its spatial variation. To avoid any ambiguity in this definition we impose additionally that $\hat{v}(k=1)=-1$, without generality loss. Under these conditions, the local dimensionless density of particles is given by $\rho\left(\bm{r}\right)=\lambda^{2}\rho\lvert\psi\left(\bm{r}\right)\rvert^{2}$, where $\rho$ represents the average particles density. It is worth noticing that within our framework the dimensionless parameter $\gamma=\lambda^{2}U\rho$ controls the relative intensity of the potential energy contribution in comparison to the kinetic energy in the ground-state energy functional, or in other words the intensity of the quantum fluctuations (zero point motion) in our system. In this way, $\gamma$ naturally represents one of the running parameter of the phase diagrams for the models considered.
To proceed, we consider that the ground-state wave function $\psi\left(\bm{r}\right)$ minimizing the energy-per-particle functional can be written in a Fourier basis. In this way, we propose [zhang2019supersolidity, mendoza2022exploring]
@@
where $c\left(\bm{k}_{j}\right)$ and $\bm{k}_{j}$ represent the Fourier amplitudes and wave vectors of the expansion. The set of wave vectors and amplitudes $\{\bm{k}_{j},c\left(\bm{k}_{j}\right)\}$ defines the type of modulated pattern under consideration and its symmetries. These wave vectors are constructed as all possible integers combinations of a vector basis $\{\bm{e}_{j}\}$, the number of elements of the basis defines the rank of the solution. In two dimensions, periodic pattern solutions correspond to configurations of rank 2 or smaller, while quasicrystals have a rank higher than 2. In Tab. 1, we present the list of solutions considered as well as its wave vector basis, here the variational parameter $k_{m}$ sets the scale of the wave vectors lattice. In terms of possible solutions, we consider all periodic patterns typically found in quasicrystal forming systems [barkan2014controlled]: one dimensional modulations, i.e., stripes, 22D hexagonal and square crystalline patterns, as well as two different kinds of compressed hexagonal patterns exciting simultaneously the two dominant characteristic wave vectors of the pair potential [lifshitz1997theoretical, jiang2016stability, yin2021transition] and finally, $n$-fold rotationally symmetric quasicrystals with $n=\{8,10,12\}$. Within the mean field spectral variational analysis, the set of Fourier amplitudes $\{c_{j}\}$ and the characteristic momentum $k_{m}$ for each possible solution are determined from the minimization of the corresponding energy-per-particle functional [prestipino2018clusterization, prestipino2018freezing]. Finally, the ground state is calculated selecting the solution with the lowest energy. For the special case of the homogeneous solution, all coefficients $c_{j}$ vanish except for $c_{0}=1$.
Pattern Vector basis $\{\bm{e}_{j}\}$ Index range
@@
Table 1: Modulated patterns considered in this work and their corresponding wave vector basis $\{\bm{e}_{j}\}$. The list of of possible solutions includes: stripes, square and hexagonal crystalline patterns, compressed hexagonal patterns exciting simultaneously wave vectors on the ratio $\tilde{k}_{2}$ (see Tab. 2) and quasicrystalline patterns with eight- (88-QC), ten- (1010-QC) or twelve-fold (1212-QC) rotational symmetry.